# Making Sense of Gravitational Thermodynamics

### v1

@@ 6 Conclusion
We have shown that equilibrium statistical mechanics and a minimal notion of thermodynamics based on coarse-grained quantities and explanations can be applied to self-gravitating systems. This analysis accounts for scientific practice and improves our understanding of the physics of SGS. More generally, our proposal sheds new light on concepts such as equilibrium and advances a new useful conceptualisation of thermodynamics standing in between statistical mechanical and phenomenological thermodynamic descriptions, which is also not agent-centric, unlike the notion of thermodynamics as control theory.
@@
Looking at future applications, this minimal conception of thermodynamics can help us understand other areas of physics, such as the physics of stars, which shares important similarities with the physics of SGS. Statistical physics can be employed to model stars, but features like stability and equilibrium work in peculiar ways and thermodynamics is usually applied only in a limited way. Another area in which statistical physics can be useful is dark matter. For instance, Chavanis et al. (2015) model dark matter halos using the King model. These systems bear interesting similarities with the SGS studied here. Finally, there are several other areas where long-range interactions are involved, including two-dimensional or geophysical fluid dynamics and plasma models.
@@ Acknowledgements
I am extremely grateful to David Wallace, Karim Thébault, and James Ladyman for their invaluable contributions and the many insightful discussions we had on the present topics, which have greatly shaped this paper. I also thank Douglas Heggie, Pierre-Henri Chavanis, and Scott Tremaine for their helpful feedback on this work. Finally, thanks to the audience at the HPS graduate students seminar at the University of Pittsburgh for their useful comments and support.

### v2

@@ 6 Conclusion and Future Directions
We have shown that equilibrium statistical mechanics and a minimal notion of thermodynamics can be applied to self-gravitating systems. Our analysis accounts for scientific practice and improves our understanding of the physics of SGS. More broadly, it sheds new light on concepts such as equilibrium and advances a new useful conceptualisation of thermodynamics in between statistical mechanical and purely phenomenological thermodynamic descriptions.
@@
Looking at future applications, this minimal conception of thermodynamics can help us understand other areas of physics, such as the physics of stars, which shares important similarities with the physics of SGS. Statistical physics can be employed to model stars, but features like stability and equilibrium work in peculiar ways. Another area in which statistical physics can be useful is dark matter. For instance, Chavanis et al. (2015) model dark matter halos using the King model. These systems bear interesting similarities with the SGS studied here. Furthermore, further areas where long-range interactions are involved include two-dimensional or geophysical fluid dynamics and plasma models. Finally, one crucial area where the results of this paper may have significant implications is the debate on black hole thermodynamics. Black holes are self-gravitating systems, and in recent years a rich body of literature has emerged questioning whether black holes can be regarded as thermodynamic objects (Dougherty and Callender (2016), Wallace (2018), Prunkl and Timpson (2019)). The debate especially focuses on the alleged analogy between the laws of black hole mechanics and the laws of phenomenological thermodynamics, though the discussion extends beyond that aspect. Black holes are sui generis entities and exhibit peculiar features that the classical SGS systems considered in this paper do not possess. Future work could apply our findings to better understand black hole thermodynamics, particularly in relation to our discussion of idealized and realistic SGS models and our proposal to introduce a minimal thermodynamic level of description. An interesting question to explore would be the extent to which black holes can be modelled using phenomenological thermodynamic descriptions, or whether they should instead be described in terms of minimal thermodynamics.
@@ Acknowledgements
This paper has greatly benefited from extensive conversations with David Wallace. I’m also grateful to Karim Thebault and James Ladyman for their valuable feedback, and to Douglas Heggie, Scott Tremaine, and Pierre-Henri Chavanis for insightful discussions. Finally, thanks to the audiences in Milan and Pittsburgh for their helpful questions and comments on earlier versions of this work. Any remaining errors are entirely my own. This paper was supported by the Swiss National Science Foundation, the SNSF Starting Grant project ‘Temporal existence’, project number TMSGI1-211294, and by the South, West and Wales doctoral training partnership.